# TUN: Detecting Significant Points in Persistence Diagrams with Deep Learning

### v1

@@ V Conclusion
-A1 Alpha Filtration
The convex hull of $n+1$ affinely independent points $\{u_{0},u_{1},\ldots,u_{n}\}\subset\mathbb{R}^{N}$ constitutes an $n$-simplex, denoted $u_{0},u_{1},\ldots,u_{n}]$. Foundational geometric primitives such as vertices, edges, triangles, and tetrahedra correspond to 0-, 1-, 2-, and 3-simplices, respectively. A simplicial complex $K$ is a finite set of simplices where two conditions hold: any face of a simplex in $K$ must also belong to $K$, and the intersection of any two simplices in $K$ is either empty or a shared face [edelsbrunner2010computational]. The dimensionality of $K$ is determined by the highest dimension of any simplex it contains. A simplicial complex $L$ is a subcomplex of $K$ if $L\subseteq K$.
@@
-A2 Persistent Homology and Persistence Diagrams
A powerful method in computational topology, persistent homology analyzes the evolution of topological features—such as connected components, tunnels, and voids—across a sequence of nested spaces [edelsbrunner2008persistent, edelsbrunner2010computational]. It quantifies the lifespan of these features, distinguishing robust structures from ephemeral noise.
@@
The output of persistent homology is commonly summarized in a persistence diagram (PD), a multiset of points in the extended plane $\mathbb{R}^{2}\cup\{\infty\}$ [edelsbrunner2010computational]. For each dimension $n$, the $n$-th PD contains a point $(b_{i},d_{i})$ for each $n$-dimensional topological feature that appears at filtration value $b_{i}$ (its birth time) and disappears at value $d_{i}$ (its death time). The vertical distance from a point to the diagonal, $d_{i}-b_{i}$, is its persistence, measuring the feature’s lifespan within the filtration.
@@
Figure 7: Some examples of point clouds.
-A3 Topology understanding of PD
@@
Now we briefly review the confidence-set method and the 2-means clustering method for topology understanding of PDs. Figure 6(a) depicts a planar point cloud whose underlying shape contains five significant loops; (b) displays the corresponding 1-PD, in which five points exhibit comparatively large persistence. For the confidence set approach (c), a statistically derived band (red) is drawn around the diagonal (with confidence level 0.5), and any point falling inside this band is regarded as noise, whereas points outside are deemed significant [fasy2014confidence]. In (d), all PD points are orthogonally projected onto the anti-diagonal line y=xy=-x and then partitioned into two clusters via 2-means. The cluster farther from the diagonal is taken as the significant set, and the closer cluster as noise.
-B Implementation Details and Hyperparameters
Input Numeric Handling
Persistence diagrams replace non-finite values by bounded sentinels. The ratio term uses $r_{i}=\log\tfrac{d_{i}}{\max(b_{i},\varepsilon)}$ with small $\varepsilon>0$ to avoid division by zero.
Persistence Diagram Encoder
Per-point descriptors augment raw $(b_{i},d_{i})$ with $p_{i}=d_{i}-b_{i}$ and $r_{i}=\log\tfrac{d_{i}}{b_{i}}$ to form $v_{i}\in\mathbb{R}^{4}$. The point-wise MLP is $4\!\rightarrow\!64\!\rightarrow\!64\!\rightarrow\!H$ with ReLU, BatchNorm and Dropout. A multi-head self-attention layer (8 heads) produces contextualized features followed by mean pooling to obtain $g_{\text{pd}}$.
Point Cloud Encoder
Operates directly on raw point clouds. The point branch applies $1\times 1$ Conv1d layers $3\!\rightarrow\!64\!\rightarrow\!64\!\rightarrow\!H$ with ReLU and BatchNorm. The global branch refines features via $H\!\rightarrow\!2H\!\rightarrow\!H$ Conv1d. Max pooling across points yields $g_{\text{pc}}$.
Multi-Modal Fusion
Linear projections transform both $g_{\text{pd}}$ and $g_{\text{pc}}$ into vectors of dimension $\tfrac{F}{2}$, and similarly, any auxiliary features $a\in\mathbb{R}^{A}$ are projected to dimension $\tfrac{F}{2}$. When auxiliary features are included, the concatenation process results in a vector $z\in\mathbb{R}^{3F/2}$. In scenarios where auxiliary features are absent, the resulting concatenated vector is $z\in\mathbb{R}^{F}$. A fusion MLP configuration of the form $D\!\rightarrow\!F\!\rightarrow\!F$, which incorporates Batch Normalization and Dropout techniques, yields the integrated context $f\in\mathbb{R}^{F}$. Here, the dimensionality $D$ is set to $3F/2$ in scenarios where auxiliary features are present; otherwise, $D$ is set equal to $F$.
Significance Classifier
For each persistence point, we concatenate its contextualized feature with $f$ and apply a per-point MLP $(H+F)\!\rightarrow\!F\!\rightarrow\!\tfrac{F}{2}\!\rightarrow\!2$ with BatchNorm and Dropout to predict logits.
Training Setup
We train with AdamW and cosine annealing; gradients are clipped to a global norm of 1.0 for stability. The batch size is 16. For each persistence diagram sample, we retain at most 100 points ($N_{\text{pd}}=100$), keeping the highest-persistence ones and padding with $(0,0)$ if fewer than 100 are present. Likewise, point clouds are capped at 50,000 points ($N_{\text{pc}}=50,000$) by random subsampling or repeated sampling padding. Here we set $N_{\text{pc}}=50,000$ because the number of points in point clouds ranges from hundreds to hundreds of thousands, and 50,000 is a relatively reasonable intermediate number.
We train with an adaptive optimizer and a monotone or annealing learning-rate schedule. To ensure stable training and prevent overfitting, we incorporate normalization across convolutional and linear layers, Dropout at multiple stages, and gradient clipping. Training stops early if validation loss does not improve for 10 consecutive epochs.
Default Configuration
Unless otherwise specified, we use $H=256$, $F=256$, attention heads $8$, and BatchNorm across hidden layers. Fusion dropout $0.3$; classifier dropouts $0.4$ then $0.3$. Data caps are $N_{\text{pd}}=100$, $N_{\text{pc}}=50,000$; batch size $16$; focal parameters $(\alpha,\gamma)=(1.0,2.0)$ with class weights $w_{0}=1.0$, $w_{1}=2.0$ for per-point loss $\ell_{i}=w_{y_{i}}\,\alpha\,\big(1-p_{t,i}\big)^{\gamma}\,\big(-\log p_{t,i}\big)$. These defaults are aligned with the released configuration and can be tuned per dataset.
-C Illustrations of Some Point Clouds
In Figure 7 we present some examples of point clouds used in our datasets, including examples of planar shapes, CAD models, TPMS models, and Zeolite structures. These examples show that the underlying models of the point clouds may illustrate complex loop features, which further show the effectiveness of our TUN model.
-D Additional Experiment Metrics and Abalation Results
TABLE III: Reports of True Positives (TP), True Negatives (TN), False Positives (FP), False Negatives (FN). As each persistence diagram sample is either truncated or padded to consist of 100 points, the total number of PD points in each dataset is $100\times\text{number of samples}$.
Planar shapes CAD models TPMS Zeolite
Model TP TN FP FN TP TN FP FN TP TN FP FN TP TN FP FN
2-means 1177 22530 1081 212 1436 45022 3119 423 4026 37128 7485 1361 4432 14321 5927 320
CS(0.5) 200 23603 8 1189 123 48138 3 1736 0 44613 0 5387 0 20248 0 4752
CS(0.9) 140 23606 5 1249 106 48138 3 1753 0 44613 0 5387 0 20248 0 4752
Ablation 1 1389 23565 46 0 1882 48098 19 1 5384 44563 50 3 4752 20163 85 0
Ablation 2 1389 23584 27 0 1859 48129 12 0 5384 44596 17 3 4752 20241 7 0
Ablation 3 1389 23602 9 0 1859 48139 2 0 5384 44599 14 3 4752 20246 2 0
Ablation 4 1389 23601 10 0 1856 48141 0 3 5382 44613 0 5 4752 20248 0 0
Ablation 5 1389 23605 6 0 1859 48140 1 0 5384 44600 13 3 4752 20246 2 0
Ablation 6 1389 23604 7 0 1859 48139 2 0 5384 44601 12 3 4752 20247 1 0
TUN 1389 23611 0 0 1883 48117 0 0 5384 44613 0 3 4752 20248 0 0

### v2

@@ III Preliminaries
III-1 Alpha Filtration
The convex hull of $n+1$ affinely independent points $\{u_{0},u_{1},\ldots,u_{n}\}\subset\mathbb{R}^{N}$ constitutes an $n$-simplex, denoted $u_{0},u_{1},\ldots,u_{n}]$. Foundational geometric primitives such as vertices, edges, triangles, and tetrahedra correspond to 0-, 1-, 2-, and 3-simplices, respectively. A simplicial complex $K$ is a finite set of simplices where two conditions hold: any face of a simplex in $K$ must also belong to $K$, and the intersection of any two simplices in $K$ is either empty or a shared face [13]. The dimensionality of $K$ is determined by the highest dimension of any simplex it contains. A simplicial complex $L$ is a subcomplex of $K$ if $L\subseteq K$.
@@ III-1 Alpha Filtration
III-2 Persistent Homology and Persistence Diagrams
A powerful method in computational topology, persistent homology analyzes the evolution of topological features—such as connected components, tunnels, and voids—across a sequence of nested spaces [12, 13]. It quantifies the lifespan of these features, distinguishing robust structures from ephemeral noise.
@@ III-2 Persistent Homology and Persistence Diagrams
The output of persistent homology is commonly summarized in a persistence diagram (PD), a multiset of points in the extended plane $\mathbb{R}^{2}\cup\{\infty\}$ [13]. For each dimension $n$, the $n$-th PD contains a point $(b_{i},d_{i})$ for each $n$-dimensional topological feature that appears at filtration value $b_{i}$ (its birth time) and disappears at value $d_{i}$ (its death time). The vertical distance from a point to the diagonal, $d_{i}-b_{i}$, is its persistence, measuring the feature’s lifespan within the filtration.
@@
III-3 Topology understanding of PD
@@
Now we briefly review the confidence-set method and the 2-means clustering method for topology understanding of PDs. Figure 4(a) depicts a planar point cloud whose underlying shape contains five significant loops; (b) displays the corresponding 1-PD, in which five points exhibit comparatively large persistence. For the confidence set approach (c), a statistically derived band (red) is drawn around the diagonal (with confidence level 0.5), and any point falling inside this band is regarded as noise, whereas points outside are deemed significant [14]. In (d), all PD points are orthogonally projected onto the anti-diagonal line y=xy=-x and then partitioned into two clusters via 2-means. The cluster farther from the diagonal is taken as the significant set, and the closer cluster as noise.
IV Methods
In this section, we introduce our novel deep learning framework, the TUN, designed for the automated classification of significance in persistence diagrams (PDs). The model leverages a multi-modal architecture that jointly processes persistence diagrams and their corresponding raw point cloud data to deliver robust and accurate predictions. Our TUN network consists of four main components: a Persistence Diagram Encoder, a Point Cloud Encoder, a Multi-Modal Fusion module, and a final Significance Classifier. The overall architecture is depicted in Figure 5. More details of implementation and hyperparameters of our TUN model are available in Appendix -A.
IV-A Input Feature Extraction and Processing
Persistence Diagram Features.
Point Cloud Data.
Auxiliary Features.
IV-B Model Architecture
Persistence Diagram Encoder.
We encode each persistence diagram as a set of enhanced per-point descriptors and aggregate them using attention [36], finally obtaining $g_{\text{pd}}$. Given raw pairs $(b_{i},d_{i})$, we construct a 4D point feature $v_{i}=[b_{i},d_{i},p_{i},r_{i}]$ with $p_{i}=d_{i}-b_{i}$ and $r_{i}=\log\tfrac{d_{i}}{b_{i}}$. Numerical stabilization is applied to avoid overflow and degeneracy. Each $v_{i}$ is processed by a point-wise MLP with standard nonlinearities and normalization, followed by a multi-head self-attention layer to model relationships among points. This yields contextualized per-point features $F_{i}$, from which a global diagram feature is obtained via pooling.
Point Cloud Encoder.
To incorporate geometric context, we use a point cloud encoder operating directly on raw point clouds, inspired by [32]. The point branch employs a stack of $1\times 1$ Conv1d layers, followed by normalization and ReLU, to generate features for each point. Subsequently, a compact global branch refines these representations, which are then aggregated into a global geometric descriptor $g_{\text{pc}}$ through a max-pooling operation.
Multi-Modal Fusion.
Significance Classifier.
IV-C Loss Function
We adopt a weighted Focal Loss to address the severe class imbalance between rare significant points and abundant noise points. For each PD point $i$, let $\mathbf{s}_{i}\in\mathbb{R}^{2}$ be the two-class logits, $\mathbf{p}_{i}=\operatorname{softmax}(\mathbf{s}_{i})$ the class probabilities with components $\mathbf{p}_{i}[0],\mathbf{p}_{i}[1]\in[0,1]$, and $y_{i}\in\{0,1\}$ the ground-truth label (1: significant, 0: noise). Denote by $p_{t,i}=\mathbf{p}_{i}[y_{i}]$ the probability assigned to the true class. The per-point focal term is $\ell_{i}=w_{y_{i}}\,\alpha\,\big(1-p_{t,i}\big)^{\gamma}\,\big(-\log p_{t,i}\big)$, where $w_{y_{i}}\in\mathbb{R}_{+}$ is the class weight of label $y_{i}$ (typically $w_{1}>w_{0}$ to upweight the positive class), $\alpha>0$ is a global scaling coefficient, and $\gamma\geq 0$ is the focusing parameter that down-weights easy examples (large $p_{t,i}$) and emphasizes hard ones. Since PDs are padded to a fixed length, we compute the loss only over valid points via a mask $M_{i}\in\{0,1\}$ and normalize by the number of valid points:
$\mathcal{L}=\frac{1}{\sum_{i}M_{i}}\sum_{i}M_{i}\,\ell_{i}.$
TABLE I: Performance comparison of our model (TUN) against baselines (Mean $\pm$ Std for machine learning methods). More details of implementation and hyperparameters of comparison models are shown in Appendix -A.
Planar shapes (250 samples) CAD models (500 samples)
Model F1 Acc Pre Rec F1 Acc Pre Rec
2-means 0.6455 0.9483 0.5213 0.8474 0.4478 0.9292 0.3153 0.7725
CS(0.5) 0.2505 0.9521 0.9615 0.1440 0.1239 0.9652 0.9762 0.0662
CS(0.9) 0.1825 0.9498 0.9655 0.1008 0.1077 0.9649 0.9725 0.0570
SVM $0.3520_{0.0277}$ $0.9561_{0.0012}$ $0.9802_{0.0017}$ $0.2148_{0.0208}$ $0.1488_{0.0094}$ $0.9619_{0.0002}$ $4363_{0.0142} 0.0$$0897_{0.0063}$
MLP $0.7536_{0.0382}$ $0.9776_{0.0027}$ $0.9677_{0.0074}$ $0.6187_{0.0535}$ $0.5743_{0.1101}$ $0.9776_{0.0044}$ $0.9534_{0.0396}$ $0.4159_{0.1075}$
[…]
TUN 0.99670.0028 0.99960.0003 0.99770.0026 0.99570.0033 0.99780.0017 0.99980.0001 0.99620.0028 0.99920.0007
TPMS (500 samples) Zeolite (250 samples)
Model F1 Acc Pre Rec F1 Acc Pre Rec
2-means 0.4765 0.8231 0.3498 0.7474 0.5866 0.7501 0.4278 0.9327
CS(0.5) 0.0000 0.8923 0.0000 0.0000 0.0000 0.8099 0.0000 0.0000
CS(0.9) 0.0000 0.8923 0.0000 0.0000 0.0000 0.8099 0.0000 0.0000
SVM $0.1212_{0.0261}$ $0.8969_{0.0009}$ $7383_{0.0202} 0.0$$663_{0.0161} 0.0$$0007_{0.0015}$ $0.8099_{0.0001}$ $1455_{0.3253} 0.0$$0003_{0.0008}$
MLP $0.7262_{0.0674}$ $0.9527_{0.0091}$ $0.9511_{0.0115}$ $0.5909_{0.0858}$ $0.6456_{0.2854}$ $0.9070_{0.0566}$ $0.9259_{0.0579}$ $0.5596_{0.3283}$
LightGBM $0.9892_{0.0079}$ $0.9976_{0.0017}$ $0.9831_{0.0157}$ $0.9955_{0.0003}$ $0.9948_{0.0057}$ $0.9980_{0.0022}$ $0.9924_{0.0114}$ $0.9974_{0.0004}$
TUN 0.99870.0007 0.99970.0002 0.99850.0009 0.99880.0008 0.99960.0004 0.99980.0002 0.99990.0001 0.99920.0008
TABLE II: Detailed performance metrics (F1-score, Accuracy, Precision, Recall) for feature ablations (Mean $\pm$ Std).
Planar shapes CAD models
Model F1 Acc Pre Rec F1 Acc Pre Rec
Ablation 1 0.98160.0144 0.99800.0016 0.98540.0118 0.97830.0267 0.96410.0391 0.99740.0029 0.97220.0442 0.95800.0548
Ablation 2 0.98900.0133 0.99880.0014 0.98700.0250 0.99120.0062 0.99520.0034 0.99960.0003 0.99150.0078 0.99670.0029
TUN 0.99670.0028 0.99960.0003 0.99770.0026 0.99570.0033 0.99780.0017 0.99980.0001 0.99620.0028 0.99920.0007
TPMS Zeolite
Model F1 Acc Pre Rec F1 Acc Pre Rec
Ablation 1 0.99610.0021 0.99910.0005 0.99540.0043 0.99680.0046 0.99440.0026 0.99790.0010 0.99020.0070 0.99860.0030
Ablation 2 0.99790.0011 0.99960.0003 0.99880.0011 0.99710.0024 0.99830.0024 0.99940.0009 0.99990.0002 0.99680.0048
TUN 0.99870.0007 0.99970.0002 0.99850.0009 0.99880.0008 0.99960.0004 0.99980.0002 0.99990.0001 0.99920.0008
Figure 6: Illustrating a failure of TUN model. In this 1-PD of TPMS test sample, three points corresponding to small-scale loops in the original model have small persistence (marked in red in Ground Truth, while blue in Prediction), which makes them hard to distinguish from noise points.
V Results
V-A Datasets
We evaluate on four categories of point-cloud data: (i) planar shapes; (ii) CAD models drawn from the ABC dataset  [^1] https://deep-geometry.github.io/abc-dataset [24] and Thingi10k dataset  [^2] https://github.com/thingi10k/thingi10k [45]; (iii) triply periodic minimal surface (TPMS) structures  [^3] https://github.com/W-W-M/Morphology-Learning [37]; and (iv) zeolite crystal frameworks  [^4] http://america.iza-structure.org/IZA-SC/ftc_table.php [1]. Our dataset comprises a total of 5,000 samples, divided into training, validation, and test sets. The training and validation sets include 3,000 samples and 500 samples, respectively. They are allocated across all four classes. The test set contains 1,500 samples divided as follows: 250 planar shape samples, 500 CAD model samples, 500 TPMS samples, and 250 zeolite samples. Each point cloud is sampled from a shape whose first Betti number is fewer than 100, guaranteeing fewer than 100 significant points in the corresponding 1-PD. Point counts per point cloud range from hundreds to 100,000, ensuring diversity in density and noise level. Each PD sample is either truncated or padded to consist of 100 points, and each point cloud is subsampled or padded to contain 50,000 points.
For each instance, we construct a point cloud by sufficiently sampling the underlying geometric model and, when specified, perturbing points with additive noise at different levels. Some examples are illustrated in Figure 7 in Appendix -B. We then compute its 1D persistence diagram. Following the established identification rules for significant points, which align with the fundamental principles of topological data analysis, significance labels are assigned by category: for planar shapes, we annotate significant points via identifying and counting loops in the planar shapes; for 3D models whose sampled surfaces form 2D manifolds (closed surfaces), we compute the genus $g$, derive the 1D Betti number $\beta_{1}$ ($\beta_{1}=2g$ for closed orientable surfaces), and label the diagram’s significant birth-death pairs in accordance with $\beta_{1}$; for surfaces with boundary or multiple connected components, we compute Betti numbers by evaluating homology group ranks based on homology theory [16] and use these to guide the labeling; and if $\beta_{1}=0$, all points are labeled as noise. Owing to sufficient sampling, significant points typically emerge at early filtration times, exhibit large persistence, and their count matches the first Betti number $\beta_{1}$ of the underlying geometric model.
Crucially, our approach avoids circularity through a distinct information asymmetry between label generation and model inference. While ground-truth labels are derived using the exact $\beta_{1}$ from the underlying mesh, the TUN model never receives $\beta_{1}$ as input. Since $\beta_{1}$ varies dynamically across samples, the model cannot simply fit a fixed persistence threshold or rely on summary statistics (which means top-$k$ persistence selection for a fixed $k$ is unfeasible for handling all samples); instead, to correctly identify the specific set of significant points, TUN is forced to leverage the geometric features from the point cloud encoder to implicitly infer the topological complexity of the shape.
V-B Evaluation Metrics
We evaluate the performance of our model using standard classification metrics, including accuracy (Acc), precision (Pre), recall (Rec), and F1-score. The required TP, TN, FP, FN metrics for computing these metrics are given in Appendix -C. The best model is selected based on the lowest loss achieved on a held-out validation set, and the main evaluation metric for testing is F1-score. This choice is motivated by the severe class imbalance in our data: for each PD, we retain the top-100 points with the largest persistence, among which the number of positive labels (significant points) is typically much smaller than that of negative labels (noise points). F1-score, as the harmonic mean of precision and recall, provides a balanced measure that is especially sensitive to the model’s ability to correctly identify the minority positive class, making it the most informative metric for this imbalanced classification task.
V-C Results and Analysis
We present the quantitative results of our TUN in Table I. The evaluation is conducted on four distinct datasets. We first compare with two classic methods: the 2-means clustering method [17] and the Confidence Set (CS) method [14]. The 2-means clustering method partitions points in the 1-PD into two clusters—significant points and noise. The confidence-region method sets a persistence threshold and labels points above it as significant. In our experiments, we used the GUDHI toolbox  [^5] https://github.com/GUDHI/TDA-tutorial/blob/master/Tuto-GUDHI-ConfRegions-PersDiag-datapoints.ipynb  [27] to compute the confidence set for each 1-PD, testing confidence levels of 0.5 and 0.9 (denoted CS(0.5) and CS(0.9)) while keeping other parameters at their defaults. Next, we compare our TUN model with three machine-learning baselines: Support Vector Machine (SVM) [19], Multilayer Perceptron (MLP) [28], and LightGBM [23]. Every machine-learning method, including ours, is run 5 times; we report the mean and standard deviation of each evaluation metric. Additionally, we also compare our TUN full model against ablation variants that rely solely on persistence diagram features or exclude the auxiliary features.
Analysis and Comparison.
Our full model, TUN, demonstrates exceptional and state-of-the-art performance across all four datasets. As shown in Table I, it consistently achieves near-perfect F1-scores with low variance across all categories. Specifically, on the four datasets, TUN achieves F1-scores of $0.9967\pm 0.0028$, $0.9978\pm 0.0017$, $0.9987\pm 0.0007$, and $0.9996\pm 0.0004$ respectively. These results establish the superiority and robustness of our deep learning-based approach, confirming that TUN can consistently and accurately identify topological features in point clouds of varying complexity. Additionally, compared with traditional and standard supervised methods, the 2-means clustering yields poor F1-scores due to low precision, while the Confidence Set method is overly conservative, leading to near-zero recall. The SVM achieves low recall (0.00030.21480.0003-0.2148), failing to capture most significant features. The MLP achieves relatively high precision, but also suffers from moderate recall (0.410.620.41-0.62), failing to capture a significant portion of topological features. The LightGBM, a gradient boosting framework, achieves competitive F1-scores across the datasets. However, statistical aggregates struggle to capture intricate topological nuances in complex geometric structures. This limitation is evident on the CAD dataset, where LightGBM’s performance drops to an F1-score of $0.9688$ with a precision of $0.9521$. Such a precision gap implies a substantial number of false positive classifications, which can be unacceptable for high-stakes topological data analysis tasks. In contrast, TUN achieves superior classification results on the same dataset. This margin highlights the necessity of our multi-modal deep learning approach: while strong baselines like LightGBM reach a performance plateau, TUN successfully bridges the final gap to reliability by learning dependencies between geometry and topology that statistical summaries fail to represent.
Ablation Studies.
We conduct ablation studies to rigorously evaluate the contribution of each component in our multi-modal framework: (1) Ablation 1: Model only has the persistence diagram encoder, without the point cloud encoder and any auxiliary features. (2) Ablation 2: TUN model without all auxiliary features, only has persistence diagram encoder and point cloud encoder. Based on Table II, the results highlight the importance of multi-modal fusion. Ablation 1 yields respectable baselines but exhibits instability, particularly on the CAD dataset (0.9641) and Planar shapes (0.9816). This suggests that relying solely on persistence values is insufficient for robust classification, as it misses crucial spatial context. Ablation 2 incorporates the point cloud encoder, leading to substantial improvements. On CAD models, F1-score jumps to 0.9952, and on Zeolite, precision reaches near perfection (0.9999). This confirms that geometric features from the raw point cloud provide essential cues that help distinguish topological features from noise, especially in complex structures. Finally, the full TUN model, which further integrates auxiliary features, achieves the best overall performance. It attains the highest F1-scores across all datasets, demonstrating that both the point cloud encoder and the auxiliary features can enhance model stability and reliability.
V-D Limitations
In this study, we restrict the number of significant points in 1-PD to fewer than 100 and the number of point cloud points to fewer than 100,000. This is due to the consideration of computational efficiency based on the assumption that the significant points are no more than 100. If the number of points in the 1-PD and the point cloud is too large, the training process will be time-consuming and memory-intensive. But theoretically, our TUN model allows handling larger datasets with larger point clouds and PD with more points. Additionally, small-scale loops are difficult for TUN to detect reliably. For example, in Figure 6, the TUN model incorrectly predicts three points that resemble noise, which in fact correspond to three small-scale loops. One possible remedy is to adopt an extremely dense sampling strategy. However, ensuring the detection of such loops via dense sampling would incur substantial computational costs. Consequently, developing methods that can accurately and efficiently capture these features remains an open challenge.
VI Conclusion
In future work, we will explore the topology understanding of higher-dimensional PDs, as well as explore alternative inputs instead of point clouds (like images, voxels) to broaden the applicability and improve the model’s interpretability.